# Molecular relaxation by reverse diffusion with time step prediction

### v1

@@ B.1 Datasets
QM7-X  [97] is a comprehensive dataset that was derived from 7000 molecular graphs sampled from the GDB13 chemical space with up to 7 heavy atoms, including types C, N, O, S, and Cl. For each SMILES string, structural and constitutional isomers were obtained using the MMFF94 force field and subsequently optimized with DFTB3+MBD computations, leading to 42 000 equilibrium structures. To capture the PES close to the equilibrium molecules, non-equilibrium molecules were generated by displacing each equilibrium molecule along a linear combination of normal mode coordinates computed with DFTB3+MBD, such that the energy difference between the non-equilibrium and equilibrium structures follow a Boltzmann distribution. For each equilibrium structure, 100 non-equilibrium configurations were generated, leading to 4 200 000 non-equilibrium structures in total, where forces and energies for each structure were computed with DFT calculations at the PBE0+MBD level with FHI-aims.
@@
We evaluate the molecular structure generation performance, discussed in details in section B.5, on the QM9 dataset [123, 124], a widely used benchmark for molecular generation tasks [82, 81, 73]. It comprises approximately 130k equilibrium organic molecules, each containing up to 9 heavy atoms of types C, O, N, and F. We use 55k molecules for training, 10k for validation, e.g. for scheduling the learning rate, and define the rest as the test split.
@@ B.3 Examples of relaxed structures
B.4 Generalization with different equivariant representations
Figure 10: The RMSD ratios (left) and RMSD (right) of molecules relaxed with the baseline MLFF model and with our MoreRed variants for $20$k non-equilibrium structures from the QM7-X test split using SO3Net implementation in SchNetPack [115] as an alternative equivariant backbone representation for PaiNN. The median values and the percentage of relaxation failure cases, i.e., the cases where the RMSD ratio exceeds $1$, are shown above each box plot.
@@ B.4 Generalization with different equivariant representations
B.5 Molecular generation with time step prediction
@@
In contrast to existing diffusion models, e.g DDPM in our case, MoreRed-AS/-JT adaptively control the reverse diffusion process with the diffusion time step predictor, as observed in the relaxation task (see Figure 3b in the main text). This means that we follow an adaptive time step schedule where the time step at each denoising/sampling iteration is estimated by a neural network $\hat{t}=\tau_{\Theta}(\mathbf{x}_{i})$. Sampling in existing diffusion models follows a pre-defined fixed schedule where exactly $T$ denoising steps with decreasing time step values $t=T,T-1,\ldots,1$ are done, as illustrated by the dashed line in the left plot in Figure 11. In the following, we compare the sampling performance with the adaptive schedule of our MoreRed to the standard, fixed diffusion model sampling process in the task of data generation from complete noise that diffusion models were originally designed for. To this end, given a molecular composition $\mathbf{Z}$, complete noise samples from the isotropic Gaussian prior distribution, $\mathcal{N}(\bm{0},\mathbf{I})$, are used as initial molecular structures and are then denoised by both models to sample valid molecular structures. For a fair comparison, we employ the same noise model for classical sampling with a fixed schedule, denoted as DDPM, and for adaptive sampling using the time step predictor, i.e. MoreRed. In the rest of this section, we refer to MoreRed-JT as MoreRed. Examples of generated structures can be found in Figure 15 and Figure 16 for MoreRed and DDPM, respectively.
We evaluate the generation performance on the QM9 dataset [123, 124], which is a widely used benchmark for molecular generation tasks and described in section B.1. For data generation with MoreRed, we set the convergence criteria to $\hat{t}\leq 0$ or a maximum number of sampling iterations equal to $2000$. For the standard DDPM [110], we use a fixed schedule with $T=1000$ sampling iterations (same as during training). For our evaluation, we generate 10 000 structures starting from the latent prior distribution where the atomic compositions $\mathbf{Z}$ of molecules are randomly drawn from the QM9 test split.
@@ C.1 Architectures
The architecture of MLFF is illustrated in Figure 12. Using the atomwise invariant features $\mathbf{s}$ from PaiNN, an atomwise multi-layer perceptron (MLP) predicts the atom-wise energies $\{E_{\mathrm{pot}}^{(i)}\}_{i=1}^{N}$, which are then aggregated to form a permutation-invariant potential energy $E_{\mathrm{pot}}=\sum_{i=1}^{N}{E_{\mathrm{pot}}^{(i)}}$, where $N$ is the number of atoms in the molecule. The gradients, i.e., the interatomic forces, are computed as the derivative of the potential energy $E_{\mathrm{pot}}$ with respect to the atom positions $\mathbf{x}$, which ensures energy conservation and equivariant predictions of interatomic forces.
@@
The architectures of the different variants are illustrated in Figure 13. On the right side is the architecture of MoreRed-JT, where the time and noise head share the same backbone molecular representation, e.g. PaiNN. Similar to the invariant energy in MLFF, the time $\hat{t}=\tau_{\Theta}(\mathbf{x}_{i})$ is predicted by an atomwise MLP using the invariant features as input. In contrast to the energy-conservative forces in MLFF, the diffusion noise is directly predicted as an equivariant tensorial quantity using an equivariant gated MLP on top of the equivariant vectorial features $\vec{\mathbf{v}}$. For MoreRed-ITP and MoreRed-AS, we use the same model architecture as for MoreRed-JT, with the only difference being that the time head $\tau_{\Theta}(\mathbf{x}_{i})$ and the noise head $\bm{\varepsilon}_{\theta}(\mathbf{x}_{i},\hat{t})$ use two separate, but identical, molecular representation networks that are trained separately, instead of using one joint network. On the left side is the architecture of the plain DDPM [110] (section 2.1) used as a baseline model in the molecular generation experiments in section B.5. The overall architecture is similar to MoreRed, except that the diffusion time step $t$ is explicitly provided by the user as input to the noise head, rather than being dynamically predicted by a neural network.

### v2

@@ B.1 Datasets
QM7-X  [104] is a comprehensive dataset that was derived from 7000 molecular graphs sampled from the GDB13 chemical space with up to 7 heavy atoms, including types C, N, O, S, and Cl. For each SMILES string, structural and constitutional isomers were obtained using the MMFF94 FF and subsequently optimized with DFTB3+MBD computations, leading to 42 000 equilibrium structures. To capture the PES close to the equilibrium molecules, non-equilibrium molecules were generated by displacing each equilibrium molecule along a linear combination of normal mode coordinates computed with DFTB3+MBD, such that the energy difference between the non-equilibrium and equilibrium structures follow a Boltzmann distribution. For each equilibrium structure, 100 non-equilibrium configurations were generated, leading to 4 200 000 non-equilibrium structures in total, where forces and energies for each structure were computed with DFT calculations at the PBE0+MBD level with FHI-aims.
@@
We evaluate the molecular structure generation performance, discussed in details in  B.6, on the QM9 dataset [130, 131], a widely used benchmark for molecular generation tasks [82, 81, 73]. It comprises approximately 130k equilibrium organic molecules, each containing up to 9 heavy atoms of types C, O, N, and F. We use 55k molecules for training, 10k for validation, e.g. for scheduling the learning rate, and define the rest as the test split.
@@ B.3 Examples of relaxed structures
B.4 Training on low energy minima from MLFF relaxations:
Figure 10: The RMSD ratios of structures relaxed with the baselines and the different MoreRed variants for 20,000 non-equilibrium structures from QM7-X, with each MoreRed variant having two versions. One version is trained on the original DFTB3+MBD equilibrium structures from QM7-X, while the other is trained on equilibrium structures resulting from further relaxation of the QM7-X reference structures using the MLFF, as in Figure 6. The median values and the percentage of failure cases, defined as those with an RMSD ratio $\geq 1$, are displayed on top of each box plot.
We investigate the structure-label mismatch in QM7-X, where force labels were computed using a more accurate and expensive method, PBE0+MBD, than for identifying the reference equilibrium structures, DFTB3+MBD. To this end, we first used the MLFF model, which was trained on the PBE0+MBD labels, to further relax the DFTB3+MBD equilibrium structures to lower energy levels. We then retrained all MoreRed variants on the relaxed data. Subsequently, we performed relaxation with the new MoreRed models on the non-equilibrium test structures from QM7-X, and computed the DFT energies of the resulting samples and their RMSD to the QM7-X reference minima, maintaining the same setup as in the main experiments.
If the inconsistency between the RMSD and energy results observed when training on the QM7-X references is due to the structure-label mismatch in the dataset, then training MoreRed on the MLFF-relaxed structures should result in similar energy and RMSD levels to those of MLFF. This is because MoreRed, by definition, learns the data manifold described by the training data and outputs structures with similar characteristics.
We summarize the results of the RMSD ratios in Figure 10 and the DFT energies in Figure 6. The three MoreRed variants achieve higher RMSD ratios but lower DFT energies, comparable to MLFF, when trained on the relaxed structures from MLFF (the lighter colors in the Figures) rather than the original DFTB3-MBD reference structures from QM7-X (the darker colors). Thus, we can conclude that the structure-label mismatch in QM7-X causes the inconsistency between RMSD and energy results observed in the original experiments. Specifically, MLFF found minima with lower energy states but different structures than the reference minima in QM7-X, resulting in higher RMSD but lower DFT energy, as it was trained on more accurate labels. These accurate labels exhibit non-negligible forces, indicating that the equilibrium structures identified by DFTB3-MBD still have higher energies.
B.5 Generalization with different equivariant representations
Figure 11: The RMSD ratios (left) and RMSD (right) of molecules relaxed with the baseline MLFF model and with our MoreRed variants for $20$k non-equilibrium structures from the QM7-X test split using SO3Net implementation in SchNetPack [122] as an alternative equivariant backbone representation for PaiNN. The median values and the percentage of relaxation failure cases, i.e. the cases where the RMSD ratio exceeds $1$, are shown above each box plot.
@@ B.5 Generalization with different equivariant representations
B.6 Molecular generation with time step prediction
@@
Figure 13: Comparison of time step trajectories for relaxation of three non-equilibrium test structures from QM7-X in red, orange, and green, similar to Figure 3b but with significantly different initial time step predictions. MoreRed-JT follows an adaptive schedule (the time step and noise are predicted at each denoising step; solid lines) and MoreRed-ITP follows a fixed schedule (the time step is only predicted at the initial denoising step; dashed lines). The dots show the predicted (but not utilized) time steps for MoreRed-ITP during denoising with the fixed schedule. The inset box shows the RMSDs between the reference equilibrium geometry and the non-equilibrium structure after and before relaxation, respectively. MoreRed-JT outperforms MoreRed-ITP in all three cases, including the orange trajectory, where MoreRed-ITP starts from a higher initial step.
In contrast to existing diffusion models, e.g DDPM in our case, MoreRed-AS/-JT adaptively control the reverse diffusion process with the diffusion time step predictor, as observed in the relaxation task (see Figure 3b in the main text for similar initial time step predictions and 13 for different initial time step predictions). This means that we follow an adaptive time step schedule where the time step at each denoising/sampling iteration is estimated by a neural network $\hat{t}=\tau_{\Theta}(\mathbf{x}_{i})$. Sampling in existing diffusion models follows a pre-defined fixed schedule where exactly $T$ denoising steps with decreasing time step values $t=T,T-1,\ldots,1$ are done, as illustrated by the dashed line in the left plot in Figure 12. In the following, we compare the sampling performance with the adaptive schedule of our MoreRed to the standard, fixed diffusion model sampling process in the task of data generation from complete noise that diffusion models were originally designed for. To this end, given a molecular composition $\mathbf{Z}$, complete noise samples from the isotropic Gaussian prior distribution, $\mathcal{N}(\bm{0},\mathbf{I})$, are used as initial molecular structures and are then denoised by both models to sample valid molecular structures. For a fair comparison, we employ the same noise model for classical sampling with a fixed schedule, denoted as DDPM, and for adaptive sampling using the time step predictor, i.e. MoreRed. In the rest of this section, we refer to MoreRed-JT as MoreRed. Examples of generated structures can be found in Figure 17 and Figure 18 for MoreRed and DDPM, respectively.
We evaluate the generation performance on the QM9 dataset [130, 131], which is a widely used benchmark for molecular generation tasks and described in B.1. For data generation with MoreRed, we set the convergence criteria to $\hat{t}\leq 0$ or a maximum number of sampling iterations equal to $2000$. For the standard DDPM [117], we use a fixed schedule with $T=1000$ sampling iterations (same as during training). For our evaluation, we generate 10 000 structures starting from the latent prior distribution where the atomic compositions $\mathbf{Z}$ of molecules are randomly drawn from the QM9 test split.
@@ C.1 Architectures
The architecture of MLFF is illustrated in Figure 14. Using the atomwise invariant features $\mathbf{s}$ from PaiNN, an atomwise multi-layer perceptron (MLP) predicts the atom-wise energies $\{E_{\mathrm{pot}}^{(i)}\}_{i=1}^{N}$, which are then aggregated to form a permutation-invariant potential energy $E_{\mathrm{pot}}=\sum_{i=1}^{N}{E_{\mathrm{pot}}^{(i)}}$, where $N$ is the number of atoms in the molecule. The gradients, i.e. the interatomic forces, are computed as the derivative of the potential energy $E_{\mathrm{pot}}$ with respect to the atom positions $\mathbf{x}$, which ensures energy conservation and equivariant predictions of interatomic forces.
@@
The architectures of the different variants are illustrated in Figure 15. On the right side is the architecture of MoreRed-JT, where the time and noise head share the same backbone molecular representation, e.g. PaiNN. Similar to the invariant energy in MLFF, the time $\hat{t}=\tau_{\Theta}(\mathbf{x}_{i})$ is predicted by an atomwise MLP using the invariant features as input. In contrast to the energy-conservative forces in MLFF, the diffusion noise is directly predicted as an equivariant tensorial quantity using an equivariant gated MLP on top of the equivariant vectorial features $\vec{\mathbf{v}}$. For MoreRed-ITP and MoreRed-AS, we use the same model architecture as for MoreRed-JT, with the only difference being that the time head $\tau_{\Theta}(\mathbf{x}_{i})$ and the noise head $\bm{\varepsilon}_{\theta}(\mathbf{x}_{i},\hat{t})$ use two separate, but identical, molecular representation networks that are trained separately, instead of using one joint network. On the left side is the architecture of the plain DDPM [117] (section 2.1) used as a baseline model in the molecular generation experiments in B.6. The overall architecture is similar to MoreRed, except that the diffusion time step $t$ is explicitly provided by the user as input to the noise head, rather than being dynamically predicted by a neural network.
@@ Appendix D Computation time
Appendix E Effect of molecular size on performance and efficiency
For many chemistry applications, it is required to perform molecular relaxation on large systems with many atoms. To evaluate how MoreRed performs on molecules with different numbers of atoms, we use the results of Section 3.2, where molecular relaxation with MoreRed-JT is performed on 20 000 non-equilibrium structures of QM7-X, and categorize the results by the number of heavy atoms per molecule. In Figure 19 we report the RMSD ratio, the number of denoising steps per relaxation trajectory and the predicted initial timestep for all structures with 5 to 7 heavy atoms. We find that, excluding some outliers, the predicted initial time step remains largely unaffected by the number of heavy atoms. Conversely, as the number of atoms grows, the median number of denoising steps needed for convergence increases by up to 10 steps, and the RMSD ratios also rise. However, using the RMSD ratios as a metric is in favor of small structures because local inaccuracies tend to affect the whole molecule, significantly impacting the atom positions at the molecule’s periphery. Therefore, similar inaccuracies at one point can lead to significantly larger RMSD of positions if the molecule is larger. Furthermore, the number of possible local minima increases with the number of heavy atoms.
Figure 19: Analysis of molecular relaxation outcomes categorized by the count of heavy atoms, spanning from 5 to 7. The graph illustrates the RMSD ratio (left), the count of denoising steps required for convergence (mid), and the estimated initial timestep (right). The initial structures are derived from the non-equilibrium configurations of QM7-X and we use MoreRed-JT for the denoising.